OpenStudy (anonymous):

how do i turn fractions to percent? like how will i know a 1/8 to percent?

OpenStudy (anonymous):

Calculate 1/8, then multiply by 100% I.e. 50/100, 50/100 = 0.50 To fraction: 0.50 x 100% = 50%

OpenStudy (anonymous):

so for 1/8 .. 12.5%?

HanAkoSolo (jamierox4ev3r):

yes, you moved the decimal point two places, which is correct

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }\times 100\]

OpenStudy (anonymous):

thanks guys :D